Prove
(a+b)/2 = a + (a-b)/2
Its basically pre-algebra but i am having a hard time wrapping my head around it this morning. i can see it is true, but cannot prove it. The reason behind it is in computer science a+b could overflow the register and provide false results. So you could do a/2 + b/2, but that uses division twice which is slower. i am still sitting here with a pen and paper trying to figure it out.

Answer by Alan3354(69443)

Prove
(a+b)/2 = a + (a-b)/2
Multiply by 2
a+b = 2a + a-b
a+b = 3a - b Not equal in general
Or -2a = -2b
a = b
True only if a = b